OpenStudy (anonymous):

Medal and FAN John has taken out a loan for college. He started paying off the loan with a first payment of $100. Each month he pays, he wants to pay back 1.1 times as the amount he paid the month before. . Explain why this series is convergent or divergent

OpenStudy (anonymous):

just tell me is this convergent or divergent

OpenStudy (solomonzelman):

well, just paying 1.1 times the previous payment means that sequence diverges (nothing to even talk about a series). BUT, a loan (I would think so) is some finite amount though, so this situtation with a loan is a bad example.....

OpenStudy (solomonzelman):

he won't go into paying more than the loan...... (and this all I am assuming there is no interest, since it is not mentioned here)

OpenStudy (solomonzelman):

So, lets write this series, and lets diregard the fact that he is to pay some finite loan amount. OK?

OpenStudy (anonymous):

is an infinite series that is not convergent, meaning that the infinite sequence of the partial sums of the series does not have a finite limit.

OpenStudy (anonymous):

is divergent

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle a_1=100 }\) \(\large\color{black}{ \displaystyle r=1.1 }\) and this is a perfect geometric series. For geometric series to converge, normally, the folloing mst be true: `|r|<1` yes correct (disregarding the fact that it's a loan).

OpenStudy (solomonzelman):

The fact that it is a loan is forcing a FINITE series.

OpenStudy (solomonzelman):

If you were to propose a different scenario, such as:

OpenStudy (solomonzelman):

John is giving a charity to poor children who stand by the Church. THe first month he gave 100$ and then, every month, he wants to give 1.1 times as much as he gave the previous month. is this series convergent or divergent?

OpenStudy (solomonzelman):

this would be more like an INFINITE series scenraio....

OpenStudy (anonymous):

ok thanks can u help with another question

OpenStudy (solomonzelman):

ok, but what course is this?

OpenStudy (solomonzelman):

is this geometry or calc?

OpenStudy (anonymous):

algebra 2

OpenStudy (solomonzelman):

oh, i see.... go ahead...

OpenStudy (solomonzelman):

((in calc there are more advanced "tools" to use))

OpenStudy (anonymous):

The graph of f(x) = 2^x + 1 is shown below. Explain how to find the average rate of change between x = 0 and x = 3

OpenStudy (solomonzelman):

find f(0). find f(3). find the slope between (0,f(0)) and (3,f(3)).

OpenStudy (solomonzelman):

making sense?

OpenStudy (anonymous):

thats the graph

OpenStudy (solomonzelman):

no need for a graph to perform this task, but, it is nice to have one.

OpenStudy (solomonzelman):

do my steps to the solution make sense?

OpenStudy (anonymous):

no

OpenStudy (solomonzelman):

Ok, lets do it togther one by one.

OpenStudy (anonymous):

so basically your finding slope

OpenStudy (solomonzelman):

yes

OpenStudy (solomonzelman):

did you find f(0) and f(3) ?

OpenStudy (solomonzelman):

f(x) = 2\(^x\)+1

OpenStudy (solomonzelman):

f(0)=2\(^0\)+1=? f(3)=2\(^3\)+1=?

OpenStudy (anonymous):

f(3) =7

OpenStudy (solomonzelman):

close

OpenStudy (solomonzelman):

2³ = ?

OpenStudy (anonymous):

9

OpenStudy (anonymous):

oops

OpenStudy (solomonzelman):

9 is f(3) or just the 2³ part?

OpenStudy (anonymous):

2^3

OpenStudy (solomonzelman):

\(2^3=2\times 2 \times 2 \)

OpenStudy (solomonzelman):

=8

OpenStudy (anonymous):

f(0) = 1

OpenStudy (anonymous):

f(3) = 9

OpenStudy (anonymous):

i meant that

OpenStudy (solomonzelman):

yes,

OpenStudy (solomonzelman):

f(3)=9

OpenStudy (solomonzelman):

and f(0)=?

OpenStudy (anonymous):

1

OpenStudy (solomonzelman):

no

OpenStudy (solomonzelman):

your function is \(f(x)=2^x+1\)

OpenStudy (solomonzelman):

\(f(0)=2^0+1=?\)

OpenStudy (anonymous):

f(0) = 0 + 1

OpenStudy (solomonzelman):

\({\rm C}^0=1\) (for any real number C)

OpenStudy (anonymous):

f(0) = 1

OpenStudy (solomonzelman):

\( f(0)=2^0+1=1+1=?\)

OpenStudy (anonymous):

2

OpenStudy (solomonzelman):

yes, f(0)=2

OpenStudy (anonymous):

Explain how to find the average rate of change between x = 0 and x = 3.

OpenStudy (solomonzelman):

now, you need to find the slope between (0,f(0)) and (3,f(3)). we know f(3)=9, f(0)=2. SO lets plug this in. ~ you have to find the slope between (0,2) and (3,9)

OpenStudy (solomonzelman):

\(\LARGE \color{black}{ \displaystyle {\rm m}=\frac{\color{blue}{{\rm y}_1}-\color{red}{{\rm y}_2}}{\color{green}{{\rm x}_1}-\color{darkgoldenrod}{{\rm x}_2}} }\) where \(\LARGE \color{black}{ \displaystyle {\rm m} }\) is the slope \(\Large\color{black}{ \displaystyle (\color{green}{{\rm x}_1}~,~~\color{blue}{{\rm y}_1}) }\) and \(\Large\color{black}{ \displaystyle (\color{darkgoldenrod}{{\rm x}_2}~,~~\color{red}{{\rm y}_2}) }\) are your two points.

OpenStudy (anonymous):

9-2 =7 3-0 =3

OpenStudy (solomonzelman):

yes, so your slope is?

OpenStudy (anonymous):

7/3

OpenStudy (solomonzelman):

yup

OpenStudy (anonymous):

is that the final answer

OpenStudy (solomonzelman):

yes, that is the final answer for your second question.

OpenStudy (anonymous):

so what is the average rate of change again

OpenStudy (solomonzelman):

7/3 :)

OpenStudy (solomonzelman):

slope and average rate of change (in this context) is the same thing.

OpenStudy (anonymous):

thanks so much

OpenStudy (solomonzelman):

you are welcome !